OpenStudy (anonymous):

Solve 3x2 = 12x. x = 0, x = 4 x = 0, x = −4 x = 3, x = 4 x = 3, x = −4

Parth (parthkohli):

Divide both sides by 3. \(\Large \color{Black}{\Rightarrow x^2 = 4x }\) Set up a quadratic equation. \(\Large \color{Black}{\Rightarrow x^2 - 4x = 0 }\) Factor. \(\Large \color{Black}{\Rightarrow x(x - 4) = 0}\) Either x = 0 -OR- x - 4 = 0 x = 4 OR x = 0

Parth (parthkohli):

x= (4,0)